Calculate the pressure, in atm, required to compress a sample of nitrogen gas from 14.55 L at 2.10 atm to 5.00 L.
Answer #1

Solution:-

From Boyle's Law,

P1V1 = P2V2

Where, P1 = initial pressure

P2 = final pressure

V1 = initial volume

V2 = final volume.

Given:- P1 = 2.10 atm , V1 = 14.55 L ; V2 = 5.00 L ; P2 = ?

( 2.10 × 14.55) = (P2 × 5.00)

or, P2 = (2.10 × 14.55) / (5.00) = 6.11 atm

hence, the pressure required = 6.11 atm

------------------------------------------------------------